# Models That Prove Their Own Correctness

### v3

@@ Appendix C Preliminaries on Stochastic Gradient Ascent
Definition C.1 (Gradient estimator).
@@
Theorem 14.8 in Shalev-Shwartz and Ben-David [2014] implies the following fact.
Fact C.2.
Fix a concave $f\colon\mathbb{R}^{d}\to[0,1]$, a norm |||||$\cdot$||| | ⋅ | | on $\mathbb{R}^{d}$, and upper-bounds $B_{\mathrm{Norm}},B_{\mathrm{Lip}}>0$. Let
$w^{\ast}\in\operatorname*{\mathrm{argmax}}_{w:||w||<B_{\mathrm{Norm}}}f(w),$
and let $\bar{w}$ denote the output of Algorithm 3 run for $N$ iterations with learning rate
@@
If at every iteration it holds that $||v_{i}||<B_{\mathrm{Lip}}$, then
@@
C.1 Learning with Stochastic Gradient Ascent/Descent
C.2 captures the general case of using SGA for maximization of concave problems. It is more common for the literature to discuss the equivalent setting of Stochastic Gradient Descent (SGD) for minimization of convex problems. Specifically, a common application of SGD is for the task of Risk Minimization: given a loss function and access to an unknown distribution of inputs, the goal is to minimize the expected loss with respect to the distribution. Assuming that the loss function is differentiable, the gradient of the loss serves as a gradient estimator (see Definition C.1) for the risk function. We refer the reader to Shalev-Shwartz and Ben-David [2014, Section 14.5.1] for a complete overview of SGD for risk minimization.
For the sake of completeness, we formulate Transcript Learning (TL, Algorithm 1) in the framework of Risk Minimization for Supervised Learning. Although multiple loss functions may achieve our ultimate goal—learning Self-Proving models—in what follows we define the loss that corresponds to TL. Fix a verifier $V$ and let $\mathcal{T}^{\ast}_{V}$ denote a distribution over accepting transcripts. We define
$\textsf{loss}\left(\theta,(x,\pi^{\ast})\right)\coloneqq\Pr_{\pi\sim\mathcal{T}^{\theta}_{V}(x)}\left[\pi\neq\pi^{\ast}\right],$ (10)
where $\pi^{\ast}$ and $\pi$ share the same verifier messages (as in Lemma B.4) so the inequality is only over the prover’s messages, namely $\Pr_{\pi\sim\mathcal{T}^{\theta}_{V}(x)}\left[\pi\neq\pi^{\ast}\right]=\Pr_{\pi\sim\mathcal{T}^{\theta}_{V}(x)}[(y,a_{1},\dots,a_{R})\neq(y^{\ast},a_{1}^{\ast},\dots,a_{R}^{\ast})]$. [^11] This loss is not to be confused with those discussed in Appendix A. Here, we are simply explaining how TL can be viewed as a supervised risk minimizer for the loss function defined in Equation 10.
@@ C.1 Learning with Stochastic Gradient Ascent/Descent
$\textsf{Risk}\left(\theta\right)\coloneqq\operatorname*{\mathbb{E}}_{\begin{subarray}{c}x\sim\mu\\ \pi^{\ast}\sim\mathcal{T}^{\ast}_{V}\end{subarray}}\left[\textsf{loss}\left(\theta,(x,\pi^{*})\right)\right],$
which means that the agreement function defined in Theorem B.5
$A(\theta)=\Pr_{\begin{subarray}{c}x\sim\mu\\ \pi^{*}\sim\mathcal{T}^{\ast}_{V}(x)\\ \pi\sim\mathcal{T}^{\theta}_{V}(x)\end{subarray}}\left[\pi=\pi^{*}\right]$
satisfies $A(\theta)=1-\textsf{Risk}(\theta)$.
Thus, maximizing the agreement is equivalent to minimizing the risk. The hypothesis class over which the optimization is performed is the ball of radius $B_{\mathrm{Norm}}$, i.e., $\left\{\theta\in\mathbb{R}^{d}:||\theta||<B_{\mathrm{Norm}}\right\}$. The assumption that $A$ is concave in $\theta$ implies that the loss function is convex in $\theta$, which is the required assumption for using SGD for risk minimization.
Indeed, TL uses the natural gradient estimator for this setting, the gradient of the “complement” of the loss: $\Pr_{\pi}\left[\pi=\pi^{\ast}\right]$, since TL maximizes the agreement instead of minimizing the risk. The proof of Lemma B.4, i.e., $\nabla_{\theta}A(\theta)=\operatorname*{\mathbb{E}}_{x,\pi^{\ast}}\left[\nabla_{\theta}\left(\Pr_{\pi}\left[\pi=\pi^{\ast}\right]\right)\right]$, follows from the above discussion.
Appendix D Annotations
We formally capture the modification described in Section 4.3 by introducing a transcript annotator and an answer extractor incorporated into the training and inference stages, respectively.
Fix a verifier $V$ in an $R$-round proof system with question length $L_{q}$ and answer length $L_{a}$. An annotation system with annotation length $\widetilde{L_{a}}$ consists of a transcript annotator $A$, and an answer extractor $E$.
In terms of efficiency, think of the annotator as an algorithm of the same computational resources as an honest prover in the system (see Definition 3.2), and the answer extractor as an extremely simple algorithm (e.g., trim a fixed amount of tokens from the annotation).
To use an annotation system the following changes need to be made:
  • At training time, an input $x$ and transcript $\pi$ is annotated to obtain $\widetilde{\pi}\coloneqq A(x,\pi)$, e.g. before the forwards backwards pass in TL (line 3 in Algorithm 1).
    The prover then extracts an answer $a_{r}\coloneqq E(\widetilde{a_{r}})$ which is sent to the verifier.
Appendix E A simple proof system for the GCD
@@
The Euclidean algorithm for computing the Greatest Common Divisor (GCD) of two integers is possibly the oldest algorithm still in use today [Knuth, 1969]. Its extended variant gives a simple proof system.
@@
Before we dive in, let us clarify what we mean by a proof system for the GCD. Prover Paul has two integers $212$ and $159$; he claims that $GCD(212,159)=53$. An inefficient way for Verifier Veronica to check Paul’s answer is by executing the Euclidean algorithm on $(212,159)$ and confirm that the output is $53$. In an efficient proof system, Veronica asks Paul for a short string $\pi^{\ast}$ (describing two integers) with which she can easily compute the answer—without having to repeat Paul’s work all over. On the other hand, if Paul were to claim that “$GCD(212,159)=51$” (it does not), then for any alleged proof $\pi$, Veronica would detect an error and reject Paul’s claim.
@@
Claim E.1 (Bézout’s identity [Bezout, 1779]).
@@
Any coefficients $z_{0},z_{1}$ satisfying the assumption of E.1 are known as Bézout coefficients for $(x_{0},x_{1})$. E.1 immediately gives our simple proof system: For input $x=(x_{0},x_{1})$ and alleged GCD $y$, the honest prover sends (alleged) Bézout coefficients $(z_{0},z_{1})$. The Verifier accepts if and only if $y=z_{0}\cdot x_{0}+z_{1}\cdot x_{1}$ and $y$ divides both $x_{0}$ and $x_{1}$.
In this proof system the Verifier does not need to make any query; to fit within Definition 3.2, we can have the verifier issue a dummy query. Furthermore, by E.1 it is complete and has soundness error $s=0$. Lastly, we note that the Verifier only needs to perform two multiplications, an addition, and two modulus operations; in that sense, verification is more efficient than computing the GCD in the Euclidean algorithm as required by Remark 3.3.
@@
To describe how a proof $z=(z_{0},z_{1})$ is annotated, let us first note how it can be computed. The Bézout coefficients can be found by an extension of the Euclidean algorithm. It is described in Algorithm 4.121212Our description follows https://en.wikipedia.org/wiki/Extended_Euclidean_algorithm.
@@
Referring to Algorithm 4, the annotation of a proof $z=(z_{0},z_{1})$ will consist of intermediate steps in its computation. Suppose that in each iteration of the While-loop, the algorithm stores each of $r_{0}$, $s_{0}$ and $q$ in an arrays $\vec{r_{0}}$, $\vec{s_{0}}$ and $\vec{q}$. The annotation $\tilde{z}$ of $z$ is obtained by concatenating each of these arrays. In practice, to avoid the transformer block (context) size from growing too large, we fix a cutoff $T$ and first trim each array to its first $T$ elements.
We formalize this in the terminology of Appendix D by defining a Transcript Annotator and Answer Extractor. Note that, since our proof system consists only of one “answer” $z$ send from the prover to the verifier, the entire transcript $\pi$ is simply $z=(z_{0},z_{1})$. Since the verification is deterministic, this means that the proof system is of an $\mathsf{NP}$ type (however, note that the search problem of finding the “$\mathsf{NP}$-witness” $z=(z_{0},z_{1})$ is in fact in $\mathsf{P}$).
  • Transcript Annotator $A$: For a fixed cutoff $T$ and given input $x=(x_{0},x_{1})$ and transcript $z=(z_{0},z_{1})$, $A$ executes Algorithm 4 on input $x=(x_{0},x_{1})$. During the execution, $A$ stores the first $T$ intermediate values of $r_{0}$, $s_{0}$ and $q$ in arrays $\vec{r_{0}}$, $\vec{s_{0}}$ and $\vec{q}$. It outputs $A(x,z)\coloneqq(\vec{r_{0}},\vec{s_{0}},\vec{q},z)$.
@@
We note that the computational complexity of $A$ is roughly that of the honest prover, i.e., Algorithm 4 (up to additional space due to storing intermediate values). As for $E$, it can be implemented in logarithmic space and linear running time in $|\tilde{z}|$, i.e., the length of the description. [^13] That is, if integers are represented by $n$-bits, then $E$ has space complexity $O(\log n+\log T)$ and running time $O(n\cdot T)$.
Appendix F Experiment details
We provide details of how we implemented the experiments in Section 5 and additional figures for each experiment. Code, data and models are available at https://github.com/orrp/self-proving-models.
@@ Appendix F Experiment details
We use Karpathy’s nanoGPT [^14] https://github.com/karpathy/nanoGPT. implementation of GPT. Note that we train the model “from scratch” only on sequences related to the GCD problem, rather than starting from a pretrained checkpoint. We use a 6.3M parameter architecture of $8$ layers, $8$ attention heads, and $256$ embedding dimensions. We optimized hyperparameters via a random hyperparameter search, arriving at learning rate $0.0007$, AdamW $\beta_{1}=0.733$ and $\beta 2=0.95$, $10\%$ learning rate decay factor, no dropout, gradient clipping at $2.0$, no warmup iterations, and $10\%$ weight decay.
@@
We sample integers from the $\log_{10}$-uniform distribution over $\{1,\dots,10^{4}\}$. Models in Tables 2 and 2 are trained for 100K iterations on a dataset of $\approx$10M samples. For Figure 3 (base ablation) we train for 20K iterations on a dataset of $\approx$1M samples; this is because this setting required 68 many runs in total, whereas the annotation-cutoff ablation required 18 longer runs.
@@
We fully describe how integer sequences are encoded. As a running example, we will use base $210$. To encode a sequence of integers, each integer is encoded in base $210$, a sign is prepended and a delimiter is appended, with a unique delimiter identifying each component of the sequence. For example, consider the input integers $x_{0}=212$ (which is $12$ in base 210) and $x_{1}=159$. Their GCD is $y=53$, with Bézout coefficients $z_{0}=1$ and $z_{1}=-1$. Therefore, the sequence $(212,159,53,1,-1)$ is encoded as
@@
As an example, consider the inputs $x_{0}=46$ and $x_{1}=39$. Tracing through the execution of Algorithm 4, we have
$\begin{array}[]{c|c|c|c|c|c|c|c|c}&x_{0}&x_{1}&y&\vec{s_{0}}&\vec{r_{0}}&\vec{q}&z_{0}&z_{1}\\ \hline\cr&46&39&&1&46&1&&\\ &&&&0&39&5&&\\ &&&&1&7&1&&\\ &&&&-5&4&1&&\\ &&&&6&3&3&&\\ &&&1&&&&-11&13\\ \end{array}$
@@
Figure 5: Verifiability as a function of the number of samples $N$. Each iteration (X axis) is a batch of 1024 samples from a dataset of $\approx$10M sequences. Every 10k iterations, Verifiability was evaluated on a held-out dataset of 1k inputs (as described in Section 5). $T$ is the number of steps in Annotated Transcript Learning (Figure 2), and $T=0$ is non-annotated Transcript Learning. Each $T$ was run with three seeds, with mean depicted by the curve and standard error by the shaded area.
Figure 6: RLVF Verifiability as a function of the number of samples $N$. Starting from a base model with Verifiability 48% (obtained via Transcript Learning), in each iteration a batch of 2048 inputs are sampled; the model generates a proof for each; the Verifier is used to check which proofs are accepted; then, the model parameters are updated accordingly (see Algorithm 2). Verifiability was evaluated on a held-out dataset of 1k inputs.

### v4

@@ A.4.1 Preliminaries on Stochastic Gradient Ascent
Definition A.3 (Gradient estimator).
@@
Theorem 14.8 in SSS14 implies the following fact.
Fact A.4.
Fix a concave $f\colon\mathbb{R}^{d}\to[0,1]$, a norm $||\cdot||$ on $\mathbb{R}^{d}$, and upper-bounds $B_{\mathrm{Norm}},B_{\mathrm{Lip}}>0$. Let
[…]
and let $\bar{w}$ denote the output of Algorithm˜3 run for $N$ iterations with learning rate
@@
If at every iteration it holds that $v_{i}||<B_{\mathrm{Lip}}$, then
@@
A.4.2 Learning with Stochastic Gradient Ascent/Descent
˜A.4 captures the general case of using SGA for maximization of concave problems. It is more common for the literature to discuss the equivalent setting of Stochastic Gradient Descent (SGD) for minimization of convex problems. Specifically, a common application of SGD is for the task of Risk Minimization: given a loss function and access to an unknown distribution of inputs, the goal is to minimize the expected loss with respect to the distribution. Assuming that the loss function is differentiable, the gradient of the loss serves as a gradient estimator (see Definition˜A.3) for the risk function. We refer the reader to SSS14 for a complete overview of SGD for risk minimization.
For the sake of completeness, we formulate Transcript Learning (TL, Algorithm˜1) in the framework of Risk Minimization for Supervised Learning. This is not strictly needed for the proof of Theorem˜A.2, but is an illuminating connection. Although multiple loss functions may achieve our ultimate goal—learning Self-Proving models—in what follows we define the loss that corresponds to TL.
Fix a verifier $V$ and let $\mathcal{T}^{\ast}_{V}$ denote a distribution over accepting transcripts. We define
$\texttt{loss}\left(\theta,(x,\pi^{\ast})\right)\coloneqq\Pr_{\pi\sim\mathcal{T}^{\theta}_{V}(x)}\left[\pi\neq\pi^{\ast}\right],$ (9)
where $\pi^{\ast}$ and $\pi$ share the same verifier messages (as in Lemma˜4.2) so the inequality is only over the prover’s messages, namely $\Pr_{\pi\sim\mathcal{T}^{\theta}_{V}(x)}\left[\pi\neq\pi^{\ast}\right]=\Pr_{\pi\sim\mathcal{T}^{\theta}_{V}(x)}[(y,a_{1},\dots,a_{R})\neq(y^{\ast},a_{1}^{\ast},\dots,a_{R}^{\ast})]$. [^9] This loss is not to be confused with those discussed in Section 3.3. Here, we are simply explaining how TL can be viewed as a supervised risk minimizer for the loss function defined in Equation 9.
@@ A.4.2 Learning with Stochastic Gradient Ascent/Descent
$\texttt{Risk}\left(\theta\right)\coloneqq\operatorname*{\mathbb{E}}_{\begin{subarray}{c}x\sim\mu\\ \pi^{\ast}\sim\mathcal{T}^{\ast}_{V}\end{subarray}}\left[\texttt{loss}\left(\theta,(x,\pi^{*})\right)\right],$
which means that the agreement function defined in Theorem˜A.2:
$A(\theta)=\Pr_{\begin{subarray}{c}x\sim\mu\\ \pi^{*}\sim\mathcal{T}^{\ast}_{V}(x)\\ \pi\sim\mathcal{T}^{\theta}_{V}(x)\end{subarray}}\left[\pi=\pi^{*}\right],$
satisfies $A(\theta)=1-\texttt{Risk}(\theta)$.
[…]
A.4.3 Proof of Theorem˜A.2
Our strategy is to cast TL as Stochastic Gradient Ascent and apply ˜A.4. Let $\varepsilon$, $B_{\mathrm{Norm}}$, $B_{\mathrm{Lip}}$ and $C$ as in the theorem statement be given. Let $\theta^{\ast}$ be such that $A(\theta^{\ast})\geq 1-\varepsilon/2$ and $\theta^{\ast}||\leq B_{\mathrm{Norm}}$.
First, notice that
$\operatorname*{\mathbb{E}}_{\bar{\theta}}\left[\mathrm{ver}_{V,\mu}(\bar{\theta})\right]\geq\operatorname*{\mathbb{E}}_{\bar{\theta}}[A(\bar{\theta})].$
This holds because, for any $x$ and model $P_{\theta}$, whenever the transcript generated by $\mathcal{T}^{\theta}(x)$ agrees with $\pi^{\ast}$, then the verifier accepts (because $\pi^{\ast}$ is honest). Therefore, to prove the theorem it suffices to show that
Following the notation in Algorithm˜1, in every iteration $i\in[N]$ the norm of the update step is
where the inequality is because $\alpha_{s}(\theta_{i})$ are probabilities, so $\leq 1$. Moreover, we have
To conclude, Lemma˜4.2 shows that TL samples from a gradient estimator for $A(\theta)$, while the above equation shows that the gradient is upper-bounded by $C\cdot B_{\mathrm{Lip}}$. We can therefore apply ˜A.4 to obtain
where the inequality is by definition of $\theta^{\ast}$. This completes the proof of Theorem˜A.2.
Appendix B Full Details of the Experimental Setup
B.1 The Bézout proof system for GCD
@@
The Euclidean algorithm for computing the Greatest Common Divisor (GCD) of two integers is possibly the oldest algorithm still in use today [Knuth69]. Its extended variant gives a simple proof system.
@@
Before we dive in, let us clarify what we mean by a proof system for the GCD. Prover Paul has two integers $212$ and $159$; he claims that $GCD(212,159)=53$. An inefficient way for Verifier Veronica to check Paul’s answer is by executing the Euclidean algorithm on $(212,159)$ and confirm that the output is $53$. In an efficient proof system, Veronica asks Paul for a short string $\pi^{\ast}$ (describing two integers) with which she can easily compute the answer—without having to repeat Paul’s work all over. On the other hand, if Paul were to claim that “GCD(212,159)=51GCD(212,159)=51” (it does not), then for any alleged proof $\pi$, Veronica would detect an error and reject Paul’s claim.
@@
Fact B.1 (Bézout’s identity [bezout]).
@@
Any coefficients $z_{0},z_{1}$ satisfying the assumption of ˜B.1 are known as Bézout coefficients for $(x_{0},x_{1})$. ˜B.1 immediately gives our simple proof system: For input $x=(x_{0},x_{1})$ and alleged GCD $y$, the honest prover sends (alleged) Bézout coefficients $(z_{0},z_{1})$. The Verifier accepts if and only if $y=z_{0}\cdot x_{0}+z_{1}\cdot x_{1}$ and $y$ divides both $x_{0}$ and $x_{1}$.
In this proof system the Verifier does not need to make any query; to fit within Definition˜3.2, we can have the verifier issue a dummy query. Furthermore, by ˜B.1 it is complete and has soundness error $s=0$. Lastly, we note that the Verifier only needs to perform two multiplications, an addition, and two modulus operations; in that sense, verification is more efficient than computing the GCD in the Euclidean algorithm as required by Remark˜3.3.
@@
To describe how a proof $z=(z_{0},z_{1})$ is annotated, let us first note how it can be computed. The Bézout coefficients can be found by an extension of the Euclidean algorithm. It is described in Algorithm˜4.101010Our description follows https://en.wikipedia.org/wiki/Extended_Euclidean_algorithm.
@@
Referring to Algorithm˜4, the annotation of a proof $z=(z_{0},z_{1})$ will consist of intermediate steps in its computation. Suppose that in each iteration of the While-loop, the algorithm stores each of $r_{0}$, $s_{0}$ and $q$ in an arrays $\vec{r_{0}}$, $\vec{s_{0}}$ and $\vec{q}$. The annotation $\tilde{z}$ of $z$ is obtained by concatenating each of these arrays. In practice, to avoid the transformer block (context) size from growing too large, we fix a cutoff $T$ and first trim each array to its first $T$ elements.
We formalize this in the terminology of Section˜6 by defining a Transcript Annotator and Answer Extractor. Note that, since our proof system consists only of one “answer” $z$ send from the prover to the verifier, the entire transcript $\pi$ is simply $z=(z_{0},z_{1})$. Since the verification is deterministic, this means that the proof system is of an $\mathsf{NP}$ type (however, note that the search problem of finding the $\mathsf{NP}$-witness” $z=(z_{0},z_{1})$ is in fact in $\mathsf{P}$).
  • Transcript Annotator $A$: For a fixed cutoff $T$ and given input $x=(x_{0},x_{1})$ and transcript $z=(z_{0},z_{1})$, $A$ executes Algorithm˜4 on input $x=(x_{0},x_{1})$. During the execution, $A$ stores the first $T$ intermediate values of $r_{0}$, $s_{0}$ and $q$ in arrays $\vec{r_{0}}$, $\vec{s_{0}}$ and $\vec{q}$. It outputs $A(x,z)\coloneqq(\vec{r_{0}},\vec{s_{0}},\vec{q},z)$.
@@
[…]
B.2 Implementation details
Code, data and models are available at https://github.com/orrp/self-proving-models.
@@ B.2 Implementation details
We use Karpathy’s nanoGPT [^12] https://github.com/karpathy/nanoGPT. implementation of GPT. Note that we train the model “from scratch” only on sequences related to the GCD problem, rather than starting from a pretrained checkpoint. We use a 6.3M parameter architecture of $8$ layers, $8$ attention heads, and $256$ embedding dimensions. We optimized hyperparameters via a random hyperparameter search, arriving at learning rate $0.0007$, AdamW $\beta_{1}=0.733$ and $\beta_{2}=0.95$, $10\%$ learning rate decay factor, no dropout, gradient clipping at $2.0$, no warmup iterations, and $10\%$ weight decay.
@@
We sample integers from the $\log_{10}$-uniform distribution over $\{1,\dots,10^{4}\}$. Models in Tables˜2 and 3 are trained for 100K iterations on a dataset of $\approx$10M samples. For Figure˜5 (base ablation) we train for 20K iterations on a dataset of $\approx$1M samples; this is because this setting required 68 many runs in total, whereas the annotation-cutoff ablation required 18 longer runs.
@@
We fully describe how integer sequences are encoded. As a running example, we will use base $210$. To encode a sequence of integers, each integer is encoded in base $210$, a sign is prepended and a delimiter is appended, with a unique delimiter identifying each component of the sequence. For example, consider the input integers $x_{0}=212$ (which is $12$ in base 210) and $x_{1}=159$. Their GCD is $y=53$, with Bézout coefficients $z_{0}=1$ and $z_{1}=-1$. Therefore, the sequence (212,159,53,1,1)(212,159,53,1,-1) is encoded as
@@
As an example, consider the inputs $x_{0}=46$ and $x_{1}=39$. Tracing through the execution of Algorithm˜4, we have
[…]